Assume that the banking system has total reserves of $200 billion. Assume also that required reserves are 20 percent of demand deposit and that banks hold no excess reserve and house hold hold no currency.

1. Find the money multiplier and money supply.
2. If the central bank now reduce required reserve to 10 percent of deposits, what will be the changes in reserve, money supply, and money multiplier?
1
Expert's answer
2014-05-20T08:22:53-0400
Total reserves = $200 billion, rr = 20%, no excess reserve, house hold no currency.
Money multiplier Mm = 1/rr = 1/0.2 = 5
Money supply = total reserves*Mm = 200*5 = $1000 billion.

2.If the central bank now reduce required reserve to 10 percent of deposits, new money multiplier will be Mm = 1/rr = 1/0.1 = 10, new total reserves will be total reserves = money supply/Mm = 1000/10 = $100 billion, money supply will be the same.